### v2

# https://vision.ist.i.kyoto-u.ac.jp/

Abstract
In this paper, we introduce a novel training method for making any monocular depth network learn absolute scale and estimate metric road-scene depth just from regular training data, i.e., driving videos. We refer to this training framework as StableCamH. The key idea is to leverage cars found on the road as sources of scale supervision but to incorporate them in the training robustly. StableCamH detects and estimates the sizes of cars in the frame and aggregates scale information extracted from them into a camera height estimate whose consistency across the entire video sequence is enforced as scale supervision. This realizes robust unsupervised training of any, otherwise scale-oblivious, monocular depth network to become not only scale-aware but also metric-accurate without the need for auxiliary sensors and extra supervision. Extensive experiments on the KITTI and Cityscapes datasets show the effectiveness of StableCamH and its state-of-the-art accuracy compared with related methods. We also show that StableCamH enables training on mixed datasets of different camera heights, which leads to larger-scale training and thus higher generalization. Metric depth reconstruction is essential in any road-scene visual modeling, and StableCamH democratizes its deployment by establishing the means to train any model as a metric depth estimator.
Keywords:
Unsupervised Monocular Depth Estimation Metric Scale Camera Height Invariance Object Size Prior
1 Introduction
Monocular depth estimation is essential for autonomous driving and advanced driver assistance systems (ADAS). It underlies many of their key perceptual tasks including 3D object detection [7, 41, 52, 34], motion planning [40, 54, 33], and road segmentation [28]. Although supervised methods have achieved impressive accuracy, they suffer from costly ground-truth data collection. This hinders their generalizability [35] and also makes continuous learning difficult.
Many recent methods [5, 55, 27, 15, 18] leverage self-supervision to avoid the need for costly supervision. Most of these methods, however, suffer from scale ambiguity which renders them difficult to use off-the-shelf. To resolve the scale ambiguity and recover the metric scale, the actual scene scale needs to be supervised in some form. Past works have adopted velocity [16], GPS [3], IMU [56, 43], or camera height [38, 48] for this. These methods require such auxiliary sensors in addition to the monocular RGB camera which significantly limits the training data. Arbitrary in-the-wild driving videos cannot be used as they lack such measurements. These self-supervised methods still require weak scale supervision and cannot be made fully self-supervised as the scale changes for each sequence captured with a different camera.
Roads, however, have plenty of objects whose projections in the video frames can inform the metric scale. Most conveniently, cars are common objects on the road on which we would like to reconstruct the depth accurately in metric scale the most. If we can somehow leverage the known sizes of cars just from their projections in the video frames, we may train a mono-depth network to be metric accurate from arbitrary driving videos. Simply learning from an object size prior would, however, be too brittle since the metric supervision will be as ambiguous as the accuracy of that prior, which likely cannot be made perfect. Not only the perspective projection, but also the variations in car makes, years, and colors, cause large variance in the size estimates, which will be directly reflected in the mono-depth errors. In addition, there are no reliable object size priors except for a few introduced as probability densities for other tasks [60, 20, 57], whose parameters are challenging to find.
How can we leverage prior knowledge about object sizes in a way robust to the actual instantiations in the scene? We derive a novel training framework which we refer to as StableCamH that answers this question. Our key idea is to aggregate object size cues in the frames and sequence into a single physical measure, the camera height. Regardless of the scene, the camera capturing it is the same for the same sequence. This means that if we can transform the object size cues in different frames into a camera height estimate, the very fact that it should not change can be used as a powerful invariance supervision to achieve metric depth reconstruction. We can formulate this as an optimization of the camera height across frames and training epochs. This leads to gradual metric-aware learning that is stable and accurate, which also makes the depth estimates consistent across the sequence.
For estimating the car sizes in each frame, we introduce Learned Size Prior (LSP). This prior estimates vehicle dimensions from its appearance and is trained with various augmentations on a large-scale dataset collected by web scraping without manual annotations. By comparing these dimension estimates in a frame with the ones computed from the estimated depth map, we obtain a per-frame scale factor. This scale can be multiplied with each camera height derived from the depth map to obtain the metric camera height for each frame. These estimated camera heights across all frames in a sequence are then consolidated by their median and used to update the supervision with weighted moving average at every end of training epochs.
We evaluate the effectiveness of StableCamH with extensive experiments using the KITTI and Cityscapes datasets and compare its accuracy with a number of weakly-supervised methods. The results clearly show that StableCamH succeeds in making monocular depth estimation methods learn metric scales and achieves the state-of-the-art accuracy despite the lack of depth and scale supervision. We also demonstrate its capability to train a model on mixed datasets collected with different cameras, i.e., camera heights. This enables larger scale training and leads to higher generalization. We believe StableCamH can serve as a fundamental building block for monocular depth estimation in the wild.
2 Related work
2.1 Self-supervision
Although, as in other tasks, supervised methods achieve the highest accuracy in monocular depth estimation (MDE) [1, 29, 24, 53, 31], they obviously require ground-truth data usually obtained with a LiDAR sensor. Self-supervised MDE methods remove the need for such hardware costs for training. Garg et al. [11] proposed the first approach to self-supervised MDE by predicting per-pixel disparities between a stereo pair. Godard et al. [14] extend this idea by introducing left-right consistency and greatly improve the accuracy. These methods require calibrated stereo data for training which limits their applicability. Zhou et al. [59] introduced a pioneering method that enables training solely from monocular videos by incorporating a network that estimates the related camera pose between successive frames. Based on this idea, many MDE methods trainable with monocular videos have been proposed [15, 27, 58, 55, 5, 18].
Despite the popularity and improvements of self-supervised MDE, directly deploying those models in autonomous driving or ADAS systems is unrealistic since they can only produce relative depth, up to an unknown scale. Our StableCamH overcomes this issue without the need for any external supervision.
2.2 Weak Supervision
Several works have tackled scale-aware MDE by adopting weak scale supervision through velocity [16], GPS [3], IMU [43], or camera height [38, 48] measurements. Zhang et al. [56] use the combination of IMU, velocity, and gravity to achieve the highest accuracy among these methods. These methods fundamentally rely on such auxiliary measurements in addition to regular RGB images. This makes it impossible to take advantage of the many road-scene videos on the Internet (e.g., YouTube) and easily installed dash cam for training. In contrast, StableCamH relies on a universal object size prior which is independent of the data collection environments and does not require any such weak supervision; it can be trained solely from monocular video sequences.
2.3 Object Size Priors
Several works for other tasks leverage priors on object sizes including single view metrology [60], object insertion [20, 57], and visual odometry [10, 36]. These methods model the priors as normal distributions or Gaussian mixture models, and the parameters of these distributions are predefined. When there is a gap between those values and the true ones, which is usually the case in the real world, the difference is directly reflected in erroneous scale estimates.
Instead of these fixed priors, we introduce a learning-based object size prior dubbed LSP. This estimates vehicle dimensions based on its appearance, thus can avoid the aforementioned problem. Furthermore, it is trained on a large-scale dataset with various augmentations, so that it can robustly estimate vehicle dimensions from any image.
3 Preliminaries
Our goal is to train a metrically accurate MDE model $\theta_{0}pt$ with only self-supervision. This self-supervision can be subdivided into two parts: one designed for learning metric scale (Sec. 4) and the other dedicated to learning depth geometry. For the latter, we follow Zhou et al. [59] and minimize the reconstruction loss between the current frame $I_{t}$ and the ones synthesized from temporally adjacent frames $I_{s},(s=t\pm 1)$. We employ a pose network $\theta_{\mathrm{pose}}$ in addition to $\theta_{0}pt$ to estimate the relative camera pose $\mathbf{T}_{t\rightarrow s}\in\mathrm{SE}(3)$ and the current dense depth map $D_{t}$, respectively:
$D_{t}=\theta_{0}pt(I_{t}),\,\,\mathbf{T}_{t\rightarrow s}=\theta_{\mathrm{pose}}(I_{t},I_{s})\,.$ (1)
We then synthesize the image from the current viewpoint $I_{s\rightarrow t}$:
$I_{s\rightarrow t}=I_{s}\expectationvalue{proj(D_{t},\mathbf{T}_{t\rightarrow s},\mathbf{K})}\,,$ (2)
where $\expectationvalue{\cdot}$ is a bilinear sampling operator, $proj(\cdot)$ outputs the 2D coordinates of $D_{t}$ when projected into the viewpoint of $I_{s}$, and $\mathbf{K}$ is the camera intrinsics. Following Godard et al. [15], we use an affine combination of $\mathrm{SSIM}$ [44] and L1 loss as the photometric error $pe$ and define the reconstruction loss $\mathcal{L}_{\mathrm{rec}}$ as the photometric error between $I_{t}$ and $I_{s\rightarrow t}$ for the pixels with minimum error across source frames:
$\mathcal{L}_{\mathrm{rec}}=\min_{s}pe(I_{t},I_{s\rightarrow t})\,,$ (3)
$pe(I_{t},I_{s\rightarrow t})=\frac{\lambda}{2}(1-\mathrm{SSIM}(I_{t},I_{s\rightarrow t}))+(1-\lambda)|I_{t}-I_{s\rightarrow t}|\,,$ (4)
where $\lambda=0.85$. We also employ an edge-aware smoothness loss [14]
$\mathcal{L}_{\mathrm{sm}}=|\partial_{x}d^{*}_{t}|e^{-|\partial_{x}I_{t}|}+|\partial_{y}d^{*}_{t}|e^{-|\partial_{y}I_{t}|}\,,$ (5)
where $d^{*}_{t}=d_{t}\big{/}\bar{d_{t}}$ is the mean-normalized inverse depth [39] to prevent the estimated depth from shrinking. Furthermore, when computing $\mathcal{L}_{\mathrm{rec}}$ we use auto-masking [15] to mask out stationary pixels, which generate incorrect supervisory signals.
Figure 1: Overview of StableCamH. At each training step $n$, an unscaled camera height $H^{\prime n,\tau}_{\mathrm{cam}}$ is differentiablly computed from the estimated depth. The previous epoch $\tau-1$ provides supervision with a scaled camera height $H^{*\tau-1}_{\mathrm{cam}}$. To compute this scaled camera height supervision, Silhouette Projector first computes the object silhouette heights $H^{\prime}_{\mathrm{obj}}$ from the depth map. By comparing $H^{\prime}_{\mathrm{obj}}$ and the estimated one $H_{\mathrm{obj}}$ with LSP, per-frame scale factor $s$ is determined and we obtain the scaled camera height $H^{n,\tau}_{\mathrm{cam}}=s\cdot H^{\prime n,\tau}_{\mathrm{cam}}$. At the end of each epoch, $H^{*}_{\mathrm{cam}}$ is optimized across a series of consecutive frames and updated with weighted moving average.
4 StableCamH
Fig. 1 depicts the overall architecture of StableCamH. Please see the caption for an overview of the framework. Let “$\,{}^{\prime}\,$” denote “not metrically scaled”.
4.1 Scale-Aware Self-Supervised Learning
In StableCamH, we derive a camera height pseudo-supervision from depth estimates. To obtain the per-frame camera height, we first compute the normal vector for each pixel $p_{i}$. We consider the 8-neighborhood of $p_{i}$ and form 8 pairs $\mathcal{N}(p_{i})=\{(p^{i}_{j_{0}},p^{i}_{j_{1}})\}^{8}_{j=1}$. Each pair $(p^{i}_{j_{0}},p^{i}_{j_{1}})$ consists of two pixels whose difference vectors w.r.t. $p_{i}$ are perpendicular to each other, and are sorted counterclockwise. We reproject the paired pixels onto 3D space and compute the cross-product of the two difference vectors w.r.t. the reprojected 3D point of $p_{i}$. By summing and normalizing all the cross-product vectors of each pair, we obtain the normal vector $\hat{\mathbf{n}}(p_{i})$
$\mathbf{n}(p_{i})=\!\!\sum_{\mathcal{N}(p_{i})}({\bf\it\phi}(p^{i}_{j_{0}})\!-{\bf\it\phi}(p_{i}))\times({\bf\it\phi}(p^{i}_{j_{1}})\!-{\bf\it\phi}(p_{i}))\,,\,\,\,\hat{\mathbf{n}}(p_{i})=\frac{\mathbf{n}(p_{i})}{\|\mathbf{n}(p_{i})\|}\,,$ (6)
where ${\bf\it\phi}(\cdot)$ denotes perspective reprojection.
For pixels in the road region, the negated inner product between the normal vector and the reprojected 3D point corresponds to the camera height from the ground plane
$H^{\prime}_{\mathrm{cam}}(p_{i})=-\,{\bf\it\phi}(p_{i})\cdot\hat{\mathbf{n}}(p_{i})\,.$ (7)
We represent the median of $H^{\prime}_{\mathrm{cam}}(p_{i})$ as the camera height of the frame
$\tilde{H}^{\prime}_{\mathrm{cam}}=\mathrm{med}(\{H^{\prime}_{\mathrm{cam}}(p_{i})\mid p_{i}\in M_{\mathrm{r}}\})\,,$ (8)
where $p_{i}$ is in the road region mask $M_{\mathrm{r}}$ obtained with an off-the-shelf semantic segmentation model [23] before training. In our training framework, the networks learn the real-world scale with the pseudo camera height, so we need the camera height in absolute scale. The camera height $\tilde{H}^{\prime}_{\mathrm{cam}}$ computed from the estimated depth map is, however, unscaled due to the lack of scale constraints from the reconstruction loss. By using Silhouette Projector, we can aggregate scale information from the object height prior into a per-frame scale factor $s$ (see Sec. 4.2) and obtain the scaled camera height $H_{\mathrm{cam}}=s\cdot\tilde{H}^{\prime}_{\mathrm{cam}}$.
The scaled camera heights $H_{\mathrm{cam}}$ computed for individual frames are inherently inaccurate due to imperfections in the estimated depth maps and the size priors. We leverage the fact that the camera height does not change in a sequence and formulate the invariance into an optimization across frames and training epochs. Thanks to this optimization, we acquire an accurate and stable supervision, which is also consistent across the scene. At the end of the $\tau$-th epoch, we use the median of $\{H^{n,\tau}_{\mathrm{cam}}\}^{N}_{n=1}$ for all frames $N$ as the representative camera height $H^{\tau}_{\mathrm{cam}}$ at the $\tau$-th epoch. Then, we update the scaled camera height pseudo-supervision $H^{*}_{\mathrm{cam}}$ with its weighted moving average:
$H^{*\tau}_{\mathrm{cam}}\leftarrow\bigg{\{}\frac{\tau(\tau-1)}{2}\cdot H^{*\tau-1}_{\mathrm{cam}}+\tau\cdot H^{\tau}_{\mathrm{cam}}\bigg{\}}\bigg{/}\frac{\tau(\tau+1)}{2}\,.$ (9)
When training on a dataset containing images captured at multiple camera heights, we create pseudo-labels corresponding to each camera height and individually optimize them.
To make the network learn the absolute scale and directly produce the scaled depth map, we impose a loss on the camera height $\mathcal{L}_{\mathrm{cam}}$ and supervise the camera height $H^{\prime}_{\mathrm{cam}}$ computed from the estimated depth map with $H^{*\tau-1}_{\mathrm{cam}}$:
$\mathcal{L}_{\mathrm{cam}}=\frac{1}{|M_{\mathrm{r}}|}\sum_{i}M_{\mathrm{r}}(p_{i})\cdot|H^{\prime}_{\mathrm{cam}}(p_{i})-H^{*\tau-1}_{\mathrm{cam}}|\,.$ (10)
Note that we use $H^{*\tau-1}_{\mathrm{cam}}$ as the pseudo-supervision during the $\tau$-th epoch training since $H^{*\tau}_{\mathrm{cam}}$ is computed at the end of this epoch.
It is worth noting that at inference we do not require either the camera height invariance assumption in a sequence or its optimization, since we employ these in order to obtain the accurate pseudo-supervision during training. The computational cost at inference is exactly the same as an original MDE model trained without StableCamH.
4.2 Silhouette Projector
In the process of creating a scaled camera height supervision, we need to know the per-frame scale factor. For this, we can consider leveraging an object of known size, i.e., we can use a prior on the object size (i.e., dimensions). To achieve this, we will need to know the 3D region which defines object dimensions. This is a non-trivial task as objects are often occluded or cropped out of images and they are in various poses.
We introduce Silhouette Projector, which measures unscaled object heights from the estimated depth map in a simple but robust manner. The scale factor can then be computed by comparing them with the height priors. The main idea of Silhouette Projector is based on two facts: 1) the height of the object silhouette projected on the plane perpendicular to the ground plane does not change regardless of the object’s pose, and 2) we can compute the silhouette height by measuring the distance from the top of the silhouette to the ground plane even if they are partly occluded or truncated, as long as the top is visible and the object is on the ground plane.
We first reconstruct an object point cloud with the estimated depth map through perspective reprojection. The object region $M_{\mathrm{obj}}$ in the image can be obtained with an off-the-shelf instance segmentation model [23] before training. We then orthographically project these 3D points onto a plane $P^{\perp}$ perpendicular to the ground plane, i.e., an arbitrary plane parallel to the road normal vector $\mathbf{\tilde{n}}$. We compute $\mathbf{\tilde{n}}$ as the median of the normal vectors $\hat{\mathbf{n}}(p_{i})$ (see Eq. 6) in the road region $M_{\mathrm{r}}$. We define the farthest distance from the 2D point on $P^{\perp}$ to the ground plane as the silhouette height. Keep in mind that this value is unscaled because we derived this from the estimated unscaled depth map. The ratio between a height prior and its silhouette height corresponds to the scale factor of interest. When there are multiple objects in an image, we use the median of the scale factors each corresponding to each object as the scale factor of the frame for computational simplicity and robustness.
4.2.1 Outlier filtering.
Some objects cannot be handled well with Silhouette Projector for various reasons including occluded upper parts, not on the ground plane, inaccurate segmented region, and misclassified object category. To filter these as outliers when computing the scale factors, we introduce an algorithm to detect them based on geometric plausibility. We first calculate the horizon $\mathbf{l}$ with the road normal vector $\mathbf{\tilde{n}}$ and the camera intrinsics $\mathbf{K}$ [17]:
$\mathbf{l}=\mathbf{K}^{-\mathsf{T}}\mathbf{\tilde{n}}\,.$ (11)
Based on the approximations in [20], we derive the approximate metric-scale object height $\hat{H}_{\mathrm{obj}}\approx\frac{h_{\mathrm{obj}}}{h_{\mathrm{cam}}}H^{*\tau-1}_{\mathrm{cam}}$ where $h_{\mathrm{cam}}$ is the farthest distance from the pixel in the object instance mask to $\mathbf{l}$, $h_{\mathrm{obj}}$ is the height of the 2D bounding box, and $H^{*\tau-1}_{\mathrm{cam}}$ is the pseudo camera height at the $\tau$-th epoch derived in Eq. 9. By comparing the relative gap between $\hat{H}_{\mathrm{obj}}$ and the object height prior $H_{\mathrm{obj}}$, we determine the outliers as
$\bigg{\{}k\,\bigg{|}\frac{|H_{\mathrm{obj}}^{k}-\hat{H}_{\mathrm{obj}}^{k}|}{H_{\mathrm{obj}}^{k}}>\mathcal{T}\bigg{\}}\,.$ (12)
In all the experiments, we set the threshold $\mathcal{T}$ to 0.2.
4.3 Training Losses
4.3.1 Auxiliary Rough Geometric Loss.
We experimentally observe that estimated object depths, especially those distant from the camera, often erroneously lean towards the road surface because of the inaccurate supervisory signal from the reconstruction loss in low-texture regions. This tendency makes silhouette heights inaccurate. To mitigate this issue, we introduce an auxiliary rough geometric loss $\mathcal{L}_{\mathrm{aux}}$, which enforces object point clouds to be within a close region. For each inlier object $k\,(k=1\dots K)$, we define $\mathcal{L}_{\mathrm{aux}}$ as the gap between the estimated depth map $D$ in the object region $M_{\mathrm{obj}}$ and the approximate object depth $D_{\mathrm{aprx}}=\frac{H_{\mathrm{obj}}}{h_{\mathrm{obj}}}f_{y}$ where $f_{y}$ is the focal length:
$\mathcal{L}_{\mathrm{aux}}\!=\!\frac{1}{K}\sum_{k}\frac{1}{|M^{k}_{\mathrm{obj}}|}\!\sum_{i}M^{k}_{\mathrm{obj}}(p_{i})\cdot\bigg{|}D(p_{i})-D^{k}_{\mathrm{aprx}}\bigg{|}\,.$ (13)
As a by-product, this loss accelerates learning the metric scale.
The $\mathcal{L}_{\mathrm{aux}}$ is potentially inaccurate since $D_{\mathrm{aprx}}$ assumes that each object can be represented by a plane parallel to the image plane, and we observe that it degrades the depth accuracy in the late stage of training. On the other hand, in the early stage, an estimated depth is unreliable, so heavily relying on the camera height loss $\mathcal{L}_{\mathrm{cam}}$ makes the training unstable. For this, we gradually decrease the weight of $\mathcal{L}_{\mathrm{aux}}$ and increase the one for $\mathcal{L}_{\mathrm{cam}}$ as the training proceeds in a logarithmic way. To ensure training stability, we stop changing these weights after the $\tau_{\mathrm{mid}}$-th epoch. This balancing can be expressed as:
$\lambda_{\mathrm{aux}}(\tau)=\begin{cases}\frac{-\log(\tau+1)}{\log(\tau_{\mathrm{mid}}+1)}&\!\!\!(\tau\leq\tau_{\mathrm{mid}})\\ \epsilon&\!\!\!(\tau>\tau_{\mathrm{mid}})\end{cases},\,\,\lambda_{\mathrm{cam}}(\tau)=\begin{cases}\frac{\log(\tau+1)}{\log(\tau_{\mathrm{mid}}+1)}&\!\!\!(\tau\leq\tau_{\mathrm{mid}})\\ 1.0&\!\!\!(\tau>\tau_{\mathrm{mid}})\end{cases},$ (14)
where $\epsilon=0.005$.
The total loss is $\mathcal{L}=\alpha\lambda_{\mathrm{cam}}\mathcal{L}_{\mathrm{cam}}+\beta\lambda_{\mathrm{aux}}\mathcal{L}_{\mathrm{aux}}+\mathcal{L}_{\mathrm{sm}}+\mathcal{L}_{\mathrm{rec}}$ where $\alpha=0.01$ and $\beta=1.0$.
5 LSP
Our StableCamH extracts metric scale from a prior on object heights through the camera height loss $\mathcal{L}_{\mathrm{cam}}$ and the auxiliary rough geometric loss $\mathcal{L}_{\mathrm{aux}}$. To ensure that StableCamH learns the scale regardless of the particular dataset used for training, a universal prior of object heights, independent of the environment, becomes essential. Cars are objects whose dimensions are explicitly and accurately known. Better yet, they are abundantly located on the road, the key semantic region whose metric depth is essential for many downstream tasks. Some methods for other tasks have modeled the size prior as a probability distribution [20, 60]. The gap between its mean and the true expectation for each environment, however, is directly reflected in the uncertainty of the scale in the supervision. Wang et al. [42] also report that there is a gap in car dimensions among datasets. These facts suggest that modeling the object size prior as a single probability distribution is problematic, especially for man-made objects like cars.
We humans not only possess rough prior knowledge about the vehicle size but can also estimate it more accurately by extracting instance-specific information such as car models from its appearance. Motivated by this intuition, we introduce a learned universal vehicle size prior termed LSP. Given a masked vehicle image, LSP estimates its dimensions.
For training, we use a dataset consisting of a large number of vehicle images and their information such as 3D dimensions. This dataset can be easily collected with web scraping and does not require manual annotations. To enhance the robustness to various input images, in addition to common color jittering and blur augmentation, we perform augmentations to simulate occlusion and truncation. For occlusion augmentation, we mask out random regions with the same shape as a vehicle or with a beam of random width. For truncation augmentation, we translate the vehicle mask in a random direction. We visualize these in supplementary material. The abundance of training images and diverse augmentations enable LSP to estimate vehicle dimensions robustly across various datasets.
Although StableCamH requires only a car height prior, we construct LSP to estimate the width and length as well as height. This enhances it to grasp an overall car size and improves the prediction accuracy of car heights.
6 Experiments
6.1 StableCamH
We evaluate the accuracy of depth estimators trained on our framework and demonstrate the effectiveness of StableCamH as a novel unsupervised metric-scale training method through depth prediction accuracy, its applicability to any network architecture, and its ability to leverage any dataset for training.
6.1.1 Implementation Details.
We implement StableCamH with PyTorch [30]. We employ Monodepth2 [15] with ResNet50 encoder pre-trained on ImageNet [32] as a default depth estimator which is one of the most basic architectures in self-supervised MDE. For all experiments unless mentioned otherwise, we train the model for 50 epochs with a batch size of 8, and set $\tau_{\mathrm{mid}}$ to 20. We use Adam optimizer [25] with an initial learning rate of $5.0\times 10^{-5}$ decayed by half every 15 epochs. During training, we perform the following augmentations with 50% chance in random order: horizontal flips, brightness adjustment ($\pm$0.2), saturation adjustment ($\pm$0.2), contrast adjustment ($\pm$0.2), and hue jitter ($\pm$0.1). For quantitative evaluation, we follow [9] and compute the seven standard metrics: AbsRel, SqRel, RMSE, $\mathrm{RMSE_{log}}$, $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$.
Table 1: Quantitative results on the KITTI Eigen test split [8] (640$\times$192 resolution). In the Supervision column, V, G, and CamH represents velocity, gravity, and camera height, respectively. The Scaling column represents whether to perform the median scaling [59]. We denote the best results in bold and the second best ones with an underline for each block. Note that we train VADepth [48] from scratch using ground-truth camera height (1.65m) provided by KITTI unlike its original paper. StabeCamH outperforms weakly-supervised methods regardless of the network architectures and without the need for ground truth supervision. In the rescaled results, every model trained with StableCamH achieves better results than its original one.
Error ($\downarrow$) Accuracy ($\uparrow$)
[…]
G2S [3] GPS - 0.109 0.860 4.855 0.198 0.865 0.954 0.980
PackNet-SfM [16] V - 0.111 0.829 4.788 0.199 0.864 0.954 0.980
Wagstaff et al. [38] CamH - 0.123 0.996 5.253 0.213 0.840 0.947 0.978
VADepth [48] CamH - 0.120 0.975 4.971 0.203 0.867 0.956 0.979
DynaDepth [56] IMU$+$V$+$G - 0.109 0.787 4.705 0.195 0.869 0.958 0.981
Ours - - 0.108 0.785 4.736 0.195 0.871 0.958 0.981
Lite-Mono [55] $+$ Ours - - 0.112 0.798 4.692 0.194 0.871 0.959 0.981
HR-Depth [27] $+$ Ours - - 0.109 0.794 4.734 0.193 0.869 0.959 0.982
VADepth [48] $+$ Ours - - 0.108 0.809 4.572 0.185 0.883 0.963 0.982
G2S [3] GPS 0.112 0.894 4.852 0.192 0.877 0.958 0.981
PackNet-SfM [16] V 0.111 0.785 4.601 0.189 0.878 0.960 0.982
Wagstaff et al. [38] CamH 0.117 0.952 4.989 0.197 0.867 0.957 0.981
VADepth [48] CamH 0.110 0.977 4.872 0.187 0.889 0.962 0.981
DynaDepth [56] IMU$+$V$+$G 0.108 0.761 4.608 0.187 0.883 0.962 0.982
Ours - 0.106 0.759 4.602 0.184 0.887 0.963 0.983
Monodepth2 [15] - 0.115 0.903 4.863 0.193 0.877 0.959 0.981
$+$Ours - 0.106 0.759 4.602 0.184 0.887 0.963 0.983
HR-Depth [27] - 0.109 0.792 4.632 0.185 0.884 0.962 0.983
$+$Ours - 0.108 0.759 4.584 0.184 0.885 0.963 0.983
Lite-Mono [55] - 0.110 0.802 4.671 0.186 0.879 0.961 0.982
$+$Ours - 0.107 0.775 4.564 0.183 0.888 0.963 0.983
VADepth [48] - 0.104 0.774 4.552 0.181 0.892 0.965 0.982
$+$Ours - 0.102 0.760 4.473 0.178 0.897 0.966 0.983
Input
Ours
Dyna
Depth [56]
G2S [3]
VADepth [48]
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionGTRefer to captionError mapRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionGTRefer to captionError mapRefer to captionRefer to captionRefer to caption
Figure 2: Qualitative comparison on KITTI. In error maps, the larger depth errors are represented in red, smaller ones in blue. The model trained with our StableCamH predicts more accurate depth map compared to the weakly-supervised methods.
6.1.2 KITTI.
Our StableCamH is independent of network architectures so we experiment on several models with StableCamH as well as Monodepth2: Lite-Mono [55] with ResNet18, HR-Depth [27] with ResNet18, and VADepth [48]. We train these models on the KITTI raw dataset [12], which contains 39,810 images for training and 4,424 for validation, and evaluate it with the KITTI Eigen test split [8] (697 images). For training and evaluation, all images are resized to 640$\times$192.
We show the qualitative results in Fig. 2. The quantitative results shown in Tab. 1 demonstrate that StableCamH enables the various architectures to learn metric scale properly. Even when employing the simplest network architecture, i.e., Monodepth2, it outperforms weakly-supervised methods that rely on ground-truth scale labels for training which our method does not need. Similar to VADepth [48], StableCamH uses the camera height invariance as supervision though VADepth requires the ground-truth and employs a complex network architecture. Nevertheless, StableCamH with Monodepth2 (row 6 in Tab. 1) surpasses it. This indicates that measuring camera height precisely is difficult and using unreliable ground-truth camera height causes low accuracy in the scale of estimated depth maps.
In the results showing the accuracy of the estimated depth map scaled with ground-truth scale factors via median scaling [59], it is evident that every model trained with StableCamH attains superior results to the one trained without StableCamH. This shows that our training framework not only enables the models to learn metric scale but also enhances the geometric accuracy. This improvement is thanks to the camera height loss $\mathcal{L}_{\mathrm{cam}}$ and $\mathcal{L}_{\mathrm{aux}}$, which makes the model respect road regions and object regions, respectively.
6.1.3 Cityscapes.
To evaluate the ability of StableCamH for taking advantage of any dataset for training, we train and evaluate it on the Cityscapes dataset [6]. We follow [59, 49, 51, 45] and collect 69,731 images for training and evaluate on the 1,525 test images. In this dataset, there are multiple camera intrinsics setups. Since this strongly affects the metric depth estimation [50], we first align the focal lengths of all images to 587.5 by resizing them, then crop them to 512$\times$192 resolution. For evaluation, we follow [2, 45] and further crop the estimated image into 416$\times$128. In order to align the number of iterations on KITTI, we train for 29 epochs with the same learning rate decreased by half every 8 epochs, and $\tau_{\mathrm{mid}}$ is set to 12.
Tab. 3 shows the quantitative results and StableCamH outperforms weakly-supervised methods by wide margins. These methods assume that highly accurate sensor data is available. Therefore, they fail to train scale-aware MDE models on datasets consisting of unreliable sensor data such as Cityscapes. On the other hand, StableCamH does not require any sensor measurement other than RGB video, so it can robustly leverage any driving videos for training.
6.1.4 Mixed datasets.
To demonstrate that StableCamH can train a model on a mixed set of datasets to attain strong generalizability, we train with 235,341 images collected from the Argoverse2 [46], Lyft [21], A2D2 [13], and DDAD [16] dataset and evaluate it on the KITTI Eigen test split [8]. Both images for training and evaluation are resized and cropped to 832$\times$512 resolution. The training images are the sets of video sequences, with each sequence being captured with its own unique camera setting. We individually optimize each camera height for each sequence during training. We detail the dataset composition and training settings in supplementary material.
As shown in Tab. 3, the model trained with StableCamH on the large-scale mixed dataset achieves comparable accuracy to the one trained on KITTI despite the domain gap and attains high generalizability. The model trained on both the mixed dataset and KITTI outperforms the model trained only on KITTI. This suggests that training on various images enhances accuracy which is the strength of fully self-supervised learning methods.
Table 2: Quantitative results on Cityscapes [6]. We train all the models from scratch with the same supervisions as the Supervision column in Tab. 1. StableCamH significantly outperforms weakly-supervised methods.
Table 3: StableCamH can be trained on mixed datasets (Argoverse2, Lyft, A2D2, and DDAD) with various camera settings, and attains high generalizability. The results are evaluated on the KITTI Eigen split [8] (832$\times$512 resolution).
Method AbsRel SqRel RMSE $\delta\!\!<\!\!1.25$
G2S [3] 0.988 23.32 28.61 0.001
PackNet-SfM [16] 0.504 6.639 14.90 0.029
VADepth [48] 0.363 7.115 11.95 0.295
StableCamH 0.125 1.288 6.359 0.858
Training dataset AbsRel SqRel RMSE $\delta\!<\!1.25$
KITTI 0.103 0.675 4.708 0.903
Mixed 0.113 0.916 5.009 0.883
Mixed $+$ KITTI 0.082 0.611 4.307 0.923
Table 3: StableCamH can be trained on mixed datasets (Argoverse2, Lyft, A2D2, and DDAD) with various camera settings, and attains high generalizability. The results are evaluated on the KITTI Eigen split [8] (832$\times$512 resolution).
6.1.5 Ablation Studies on KITTI.
To study the effect of the proposed components in StableCamH, we conduct ablation studies on the training losses, balancing the loss weights, camera height optimization, and outlier filtering. The results are shown in Tab. 5. Thanks to the accuracy of LSP and the robustness of Silhouette Projector, the camera height loss makes a greater contribution to the depth accuracy than the auxiliary rough geometric loss. By introducing the camera height optimization, the model can learn the scale with more accurate and consistent supervision than leveraging the size prior independently for each frame (rows 1, 2, and 6 in Tab. 5), which leads to improved accuracy. Though balancing the loss weights (Eq. 14) with either $\mathcal{L}_{\mathrm{aux}}$ or $\mathcal{L}_{\mathrm{cam}}$ degrades the results due to the insufficient strength of the supervisory signals for metric scale, it contributes to improving accuracy when both $\mathcal{L}_{\mathrm{aux}}$ and $\mathcal{L}_{\mathrm{cam}}$ are employed.
Our proposed method computing the camera height pseudo-supervision from estimated depth maps with Silhouette Projector and LSP can also be used with a pre-trained MDE model in an offline manner. We first train the original Monodepth2 [15] without StableCamH, then predict depth maps of training images. With Silhouette Projector and LSP, we obtain per-frame scaled camera height and optimize it across frames with the median operation. We fix this value and use it as the camera height supervision for an untrained Monodepth2 on StableCamH. We examine the difference between the proposed camera height optimized online and the fixed one derived offline when we use them as the pseudo-supervision. Note that we also fix $\lambda_{\mathrm{cam}}$ because the pseudo-label can be considered sufficiently accurate from the early stage of training. We also run another experiment in which we fix camera height supervision derived offline for the first 20 epochs, which is the same number of training epochs as the original Monodepth2, then unfix the camera height pseudo-label to fine-tune it via online optimization. As shown in Tab. 5, the model supervised with the fixed camera height computed offline achieves slightly better results than the one supervised with the camera height optimized online. This result indicates that our strategy of deriving scaled camera height is highly accurate and fixing it stabilizes the training. The model trained with the fine-tuned camera height achieves the highest accuracy, which means our StableCamH can attain better accuracy with additional pre-training costs.
Table 4: Each loss contributes to learning metric scale and the others enhance the depth accuracy. Bal: balancing the loss weights (Eq. 14). Opt: camera height optimization. Flt: outlier filtering.
Table 5: Training with a fine-tuned pseudo camera height achieves the highest accuracy. Online: optimizing the camera height while training. Offline: training with a fixed one computed from depth maps predicted with pre-trained Monodepth2. Fine-tune: training with the fixed one computed offline for 20 epochs, then optimizing online.
$\mathcal{L}_{\mathrm{aux}}$ $\mathcal{L}_{\mathrm{cam}}$ Bal Opt Flt AbsRel SqRel RMSE $\delta\!\!<\!\!1.25$
- - - 0.125 0.887 5.055 0.834
- - 0.149 0.950 5.219 0.785
- - 0.116 0.875 4.874 0.863
- 0.211 1.303 6.143 0.585
- 0.122 0.880 5.017 0.849
- 0.115 0.822 4.811 0.858
- 0.116 0.823 4.901 0.860
0.108 0.785  4.736 0.871
Deriv. method AbsRel SqRel RMSE $\delta\!<\!1.25$
Online 0.108 0.808 4.740 0.870
Offline 0.106 0.790 4.719 0.873
Fine-tune 0.105 0.782 4.682 0.876
Table 5: Training with a fine-tuned pseudo camera height achieves the highest accuracy. Online: optimizing the camera height while training. Offline: training with a fixed one computed from depth maps predicted with pre-trained Monodepth2. Fine-tune: training with the fixed one computed offline for 20 epochs, then optimizing online.
6.2 LSP
6.2.1 Implementation Details.
We implement LSP with PyTorch [30]. We train it for 15 epochs with a batch size of 32 and warm-up the first 5 epochs. We use AdamW [26] optimizer with a learning rate of $5.0\times 10^{-4}$ linearly decreased to $4.0\times 10^{-4}$ and a weight decay of 0.05. As a backbone network, we employ ConvNeXt V2 [47] pre-trained with ImageNet22K [32] and concatenate two linear layers to it. As a training dataset, we use DVM-CAR [22], which contains 1,451,784 vehicle mask images from 899 UK market car models and their dimensions. All images are resized to 300$\times$300 resolution. We use 80% of the data for training and others for validation. We adopt multiple data augmentations mentioned in Sec. 5 to make the model robust to the domain gap. We use L1 loss for each dimension as the objective function.
6.2.2 Evaluation on KITTI and Cityscapes.
We evaluate the generalizability of our LSP with the dimensions of the annotated 3D bounding box on both the KITTI [12] and the Cityscapes [6] datasets in the Car class. We compare the results of LSP to a fixed car height prior (1.59m, following [20]) and MonoCInIS [19], a 3D object detection model that outputs object dimensions from an image and an instance mask. We train MonoCInIS from scratch with the KITTI dataset. As an evaluation metric, we compute the absolute relative error for each dimension. When we estimate car dimensions with LSP and MonoCInIS, we use a ground-truth instance mask for making a car image mask and as input, respectively. For all the experiments in LSP, input vehicle masks are cropped and resized to 300$\times$300 resolution. On the KITTI dataset, we extract 3,769 samples for evaluation [4]. On the Cityscapes dataset, we evaluate on the validation split, which has 500 samples. On the KITTI dataset, each object is separated into 3 levels based on detection difficulty: Easy, Moderate, and Hard. Considering the difference in focal lengths between KITTI and Cityscapes, we also separate objects on Cityscapes into the same levels as on KITTI.
As shown in Tab. 6, LSP outperforms the other two methods in most cases. Particularly, the accuracy of MonoCInIS drops on Cityscapes since it overfits the training dataset, i.e., KITTI and cannot address the domain gap. In contrast, thanks to the large training data and various augmentations, our LSP obtains the highest generalizability and robustly estimates dimensions on any dataset. Learning not only car height but also width and length also improves the prediction accuracy as this facilitates to understand overall car sizes.
Table 6: Generalizability on KITTI and Cityscapes. We evaluate LSP on Absolute Relative Error for each dimension (height/width/length), comparing with the fixed car height prior (1.59 m, following [20]) and MonoCInIS [19] trained on KITTI. LSP robustly surpass other methods in most cases and can predict accurately regardless of the datasets.
Method Easy Moderate Hard All
KITTI
Fixed Height [20] 0.065 / - / - 0.064 / - / - 0.065 / - / - 0.065 / - / -
MonoCInIS [19] 0.069 / 0.065 / 0.093 0.112 / 0.106 / 0.133 0.268 / 0.263 / 0.269 0.159 / 0.155 / 0.177
LSP (only height) 0.057 / - / - 0.089 / - / - 0.106 / - / - 0.087 /
LSP 0.046 / 0.136 / 0.130 0.060 / 0.144 / 0.139 0.073 / 0.155 / 0.150 0.062 / 0.147 / 0.145
Cityscapes
Fixed Height [20] 0.093 / - / - 0.093 / - / - 0.076 / - / - 0.092 / - / -
MonoCInIS [19] 0.578 / 0.599 / 0.584 0.759 / 0.766 / 0.767 0.843 / 0.852 / 0.852 0.851 / 0.859 / 0.856
LSP (only height) 0.063 / - / - 0.075 / - / - 0.103 / - / - 0.096 /
LSP 0.056 / 0.069 / 0.067 0.068 / 0.068 / 0.078 0.069 / 0.059 / 0.075 0.080 / 0.064 / 0.077
7 Conclusion
In this work, we proposed StableCamH, a novel training framework to make any MDE model learn metric scale. StableCamH aggregates the scale information obtained from a learned object size prior into the camera height and uses it as pseudo-supervision formulated as optimization across frames and epochs by leveraging its invariance. We showed that StableCamH enables monocular depth estimators to grasp metric scale regardless of network architectures and achieves state-of-the-art absolute-scale depth estimation accuracy on both the KITTI and the Cityscapes dataset, and demonstrated that it can leverage any dataset for training. We also proposed LSP, which can be easily trained on a large-scale dataset without manual annotations and robustly estimates vehicle dimensions from their appearances. We showed that the learned size prior is superior to a fixed one and outperforms a similar 3D object detection model. StableCamH can be leveraged as a drop-in training scheme for any monocular depth estimation method. We hope it will be a versatile tool for road scene understanding research.
Acknowledgements
This work was in part supported by JSPS 20H05951 and 21H04893, JST JPMJCR20G7 and JPMJAP2305, and RIKEN GRP.
Appendix 0.A Details of Training on the Mixed Datasets
0.A.1 Dataset Preparation
We use the images from the front camera in Argoverse2 [46], Lyft [21], A2D2 [13], and DDAD [16] as training datasets. As mentioned in Sec. 6.1 in the main paper since the focal lengths are essential for metric MDE, we resize the image and align the focal lengths to 1000 px. In order to facilitate the training process, we crop the images to 832$\times$512 resolution. These resizing and cropping are also conducted during evaluation on KITTI Eigen split [8].
To make the training easier and more efficient, we eliminate the frames where the camera is stationary, since training with these frames greatly degrades the inference performance [15]. We achieve this elimination by looking at pairs of successive frames and simply checking the number of pixels with large differences in normalized intensities. Images with a small number of pixels with large intensity changes are deemed stationary and discarded from the training dataset. Since Argoverse2 is captured at a higher frame rate than the other datasets, we treat every other frame as successive frames.
For optimization during training, StableCamH requires a sufficient number of images captured with the same camera height. We select such sequences that include at least 2,500 frames. At every end of the training epoch, we optimize each camera height individually. In total, we use 235,341 images for training.
0.A.2 Training Settings
We train Monodepth2 [15] with ResNet50 on StableCamH with the same batch size and initial learning rate as used during training on the KITTI dataset (Sec. 6.1 in the main paper). Considering the gap of training samples between KITTI and the mixed datasets, we set $\tau_{\mathrm{mid}}=4$ and train the model for 8 epochs while reducing the learning rate by half every 3 epochs.
Appendix 0.B Ablation of LSP
Figure 3: Visualization of beam/instance occlusion and truncation augmentations for training LSP.
As specified in Sec. 5 in the main paper, during training of LSP, we perform beam/instance occlusion and truncation augmentations illustrated in  Fig. 3 along with common color jittering and blur. The beam occlusion simulates objects occluded by poles including traffic lights and trees. Their segmentation mask is divided into more than two regions. The instance occlusion augmentation is for objects occluded by the other stuffs including vehicles and people. The truncation one is for objects that are partially cropped out of images. In  Tab. 7 we ablate the effectiveness of these augmentations. The results show that each augmentation contributes to improving the prediction accuracy both on the KITTI and Cityscapes datasets. This indicates that they enhance the robustness to any input image.
Table 7: Ablation of augmentations for LSP on KITTI [4] and Cityscapes [6]. We evaluate each predicted dimension (height/width/length) on the absolute relative error. Both beam/instance occlusion and truncation augmentation improve the prediction accuracy.
beam occ. inst. occ. trunc. KITTI Cityscapes
0.081 / 0.159 / 0.174 0.088 / 0.065 / 0.082
0.071 / 0.162 / 0.179 0.080 / 0.069 / 0.080
0.074 / 0.150 / 0.172 0.083 / 0.065 / 0.078
0.079 / 0.136 / 0.165 0.084 / 0.057 / 0.077
0.062 / 0.147 / 0.145 0.080 / 0.064 / 0.077
Appendix 0.C Additional Quantitative Results on KITTI
In the main paper, we trained the multiple networks with StableCamH on the KITTI dataset [12] and evaluated them on the KITTI Eigen split [8] which is the most widely used, comparing the weakly-supervised methods. We also evaluate the same models with the improved ground-truth depth maps for KITTI introduced in [37]. These improved depth maps are composed of 652 out of 697 test frames in the Eigen split. The results shown in Tab. 8 demonstrate that the models trained with StableCamH still surpass weakly-supervised methods, both with and without the median scaling, as well as when evaluating on the KITTI Eigen split. It should be emphasized that VADepth [48] trained with StableCamH (row 9 and 18) outperforms the same network supervised with the measured camera height (row 4 and 13). This proves that the pseudo camera height obtained via the camera height optimization is more precise than the measured one.
Table 8: Evaluation on KITTI improved ground truth [37] (640$\times$192 resolution). The models trained with StableCamH outperform weakly-supervised methods in both results, with and without the median scaling [59].
Error ($\downarrow$) Accuracy ($\uparrow$)
[…]
G2S [3] GPS - 0.111 0.791 4.523 0.168 0.869 0.967 0.989
PackNet-SfM [16] V - 0.105 0.656 4.098 0.156 0.886 0.971 0.991
Wagstaff et al.  [38] CamH - 0.097 0.636 4.281 0.153 0.887 0.975 0.993
VADepth [48] CamH - 0.091 0.555 3.871 0.134 0.913 0.983 0.995
DynaDepth [56] IMU + V + G - 0.092 0.477 3.756 0.135 0.906 0.983 0.996
Ours - - 0.089 0.506 3.806 0.134 0.908 0.984 0.996
Lite-Mono [55] $+$ Ours - - 0.090 0.526 3.767 0.134 0.910 0.984 0.995
HR-Depth [27] $+$ Ours - - 0.096 0.524 3.852 0.139 0.902 0.982 0.996
VADepth [48] $+$ Ours - - 0.101 0.547 3.675 0.133 0.910 0.984 0.996
G2S [3] GPS 0.088 0.544 3.968 0.137 0.913 0.981 0.995
PackNet-SfM [16] V 0.100 0.606 3.943 0.145 0.900 0.977 0.993
Wagstaff et al.  [38] CamH 0.095 0.602 4.145 0.146 0.902 0.978 0.994
VADepth [48] CamH 0.080 0.518 3.745 0.123 0.933 0.986 0.996
DynaDepth [56] IMU + V + G 0.084 0.448 3.761 0.130 0.917 0.984 0.996
Ours - 0.081 0.455 3.675 0.125 0.927 0.986 0.996
Lite-Mono [55] $+$ Ours - 0.083 0.478 3.632 0.126 0.926 0.986 0.996
HR-Depth [27] $+$ Ours - 0.084 0.461 3.719 0.129 0.922 0.985 0.996
VADepth [48] $+$ Ours - 0.077 0.426 3.455 0.116 0.937 0.989 0.997
Input
Ours
Dyna
Depth [56]
G2S [3]
VADepth [48]
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionGTRefer to captionError mapRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionGTRefer to captionError mapRefer to captionRefer to captionRefer to caption
Input
Ours
Dyna
Depth [56]
G2S [3]
VADepth [48]
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionGTRefer to captionError mapRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionGTRefer to captionError mapRefer to captionRefer to captionRefer to caption
Figure 4: Additional qualitative comparison on KITTI. We compare estimated depth maps of Monodepth2 [15] trained with StableCamH (Ours) with the ones of weakly-supervised methods: G2S [3], DynaDepth [56], and VADepth [48]. In error maps, the larger depth errors are represented in red, and smaller ones are depicted in blue. The results show that StableCamH can predict more accurate depth maps in various scenes, compared with the weakly-supervised methods.
Appendix 0.D Additional Qualitative Results on KITTI
We provide additional results of StableCamH in diverse scenes on the KITTI dataset. As shown in Fig. 4, StableCamH can predict more accurate depth maps than the weakly-supervised methods, especially in road and object regions.